If n(U)=24,n(A)=11,n(B)=18,find the greatest number of n(AUB)?
Answer by stanbon(75887)

If n(U)=24,n(A)=11,n(B)=18,find the greatest number of n(AUB)?
---------
n(AUB) = n(A)+n(B)-n(A and B)
-----
n(AUB) = 11 + 18 - n(A and B) <= 24
29 - n(A and B) <= 24
n(A and B) >= 5
-----
If n(A and B) = 5, n(AUB) = 24
-----------------------
Cheers,
Stan H.
---------------